# IDEA: An Inverse Domain Expert Adaptation Based Active DNN IP Protection Method

### v1

@@ III-C SteganoGAN
SteganoGAN [33] is a lightweight steganographic network for key encoding and decoding. It consists of three sub-networks: the encoder $\mathcal{E}(\cdot,\cdot)$, the decoder $\mathcal{D}(\cdot)$, and the critic $\mathcal{C}(\cdot)$. The training loss of the encoder-decoder network is composed of three losses, $\mathcal{L}_{d}$, $\mathcal{L}_{s}$ and $\mathcal{L}_{r}$:
$\mathcal{L}_{d}=\mathbb{E}_{x\sim\mathcal{P}_{C}}\left[\mathcal{L}_{\text{BCE}}\left(\mathcal{D}\left(\mathcal{E}(x,\emph{msg}),\emph{msg}\right)\right)\right],$ (9)
$\mathcal{L}_{s}=\mathbb{E}_{x\sim\mathcal{P}_{C}}\left[\mathcal{L}_{\text{MSE}}\left(x-\mathcal{E}(x,\emph{msg})\right)\right],$ (10)
[…]
where $\mathcal{P}_{C}$ denotes the training set of SteganoGAN, $\mathcal{L}_{\text{BCE}}$ and $\mathcal{L}_{\text{MSE}}$ represent the binary cross-entropy loss and mean square error loss, respectively. msg is an encoded or a decoded binary message that has the same dimension as the cover image $x$. The decoding loss $\mathcal{L}_{d}$ ensures that the encoded message can be accurately extracted from the stego image; the similarity loss $\mathcal{L}_{s}$ forces the stego image to resemble the cover image; and the realness loss $\mathcal{L}_{r}$ mitigates distortion in the stego image.
@@ III-D Domain Expected Risk Bound
$\mathcal{R}_{\mathcal{T}}(f)\leq$ $\mathcal{R}_{\mathcal{S}}(f)-4I(z;\hat{z})$ (12)
$+4H(Y)+\frac{1}{2}d_{\mathcal{H}\Delta\mathcal{H}}\left(p(z),p(\hat{z})\right),$
@@
The unknown conditional distribution $p(\hat{z}|z)$ can be approximated by the variational distribution $q_{\phi}(\hat{z}|z)$ with an auxiliary neural network parameterized with $\phi$ [36]. By feeding samples $\{(z_{i},\hat{z}_{i})\}_{i=1}^{N_{\text{est}}}$ into the auxiliary network,
@@
Figure 1: Overall pipeline of IDEA, a solution to the inverse problem of domain adaptation for key-based active authorization.
A DNN classifier $f$ of $L$ layers can be expressed as $f=f_{1}\circ f_{2}\circ\cdots\circ f_{L}$. The activation output of the $l$-th layer is denoted as $f^{l}\in\mathbb{R}^{{c_{l}}\times{h_{l}}\times{w_{l}}}$, where $c_{l}$, $h_{l}$ and $w_{l}$ denote the dimensions of channel, height and width, respectively of the activation map. Previous works [35, 26] focused on minimizing the MI between the latent features output by the penultimate layer (i.e., $f_{L-1}$) to reduce the generalization of $\mathcal{T}$. We observe that the source and target domain samples share minimal information at the penultimate layer. Their latent features in the earlier layers are fairly similar, which can leak vital information about the protected model’s functionality. To avoid this vulnerability, we jointly minimize the MI between $\mathcal{S}$ and $\mathcal{T}$ on 4 selected hidden layers that are evenly distributed from shallow to deep within the model, i.e.,
$\mathop{\min}\sum_{l\in T_{\text{sel}}}\hat{I}(f^{l};\hat{f}^{l}),\quad\text{where }T_{\text{sel}}=\{l_{1},l_{2},l_{3},l_{4}\}.$ (15)
@@ IV-A Key Embedding
Figure 2: Key embedding into a cover image by the pretrained SteganoGAN encoder and key extraction from the stego image by the model owner with the corresponding pretrained SteganoGAN decoder (private). The SteganoGAN encoder and decoder are pre-trained on the original training dataset $D_{\text{tr}}$.
Fig. 2 shows an example of key encoding and decoding. Each individual user is assigned a unique key ${k_{j}}\in{\{0,1\}}^{{c}\times{r}\times{r}}$, which is a binary square matrix derived from the identity of a user, where $c$ is the number of channels and $r$ denotes the length of the square matrix. The entire key space $\mathcal{K}$ is $2^{{c}\times{r}\times{r}}$. Notably, $r$ is much smaller than the height $h$ and width $w$ of the original image $x\in{[0,1]}^{{3}\times{h}\times{w}}$. To match the sizes of $k_{j}$ and $x$, we concatenate multiple non-overlapping blocks of $k_{j}$ to obtain the expanded key $\text{Exp}(k_{j})\in{\{0,1\}}^{{c}\times{h}\times{w}}$. The encoded image $x_{j}=\mathcal{E}(x,\text{Exp}(k_{j}))$ is generated by simultaneously feeding $x_{i}$ and $\text{Exp}(k_{j})$ into the encoder. During the verification stage, $x_{j}$ is passed through the decoder $\mathcal{D}(\cdot)$ and split into multiple non-overlapping blocks of size ${r}\times{r}$. A majority voting process is then conducted across corresponding elements of all blocks to extract the key $\hat{k}_{j}$. This operation can be formally expressed as $\hat{k}_{j}=\text{Vote}(\text{Split}(\mathcal{D}(x_{j})))$.
The encoder is distributed to all authorized users for embedding their unique user keys into the test images. The decoder is kept private and reserved for the model owner to recover the embedded user key in any encoded test images. The perturbations introduced by the encoder into the images are visually imperceptible and sample-specific, i.e., the perturbations due to the same embedded key vary across images.
@@ IV-B Experts Training
Let $\mathcal{A}_{j}$ be the source domain, and $\mathcal{B}$ and $\mathcal{N}_{j}$ be two different target domains. An inverse problem of domain adaption can be solved by optimizing the domain expected risk bound in Sec. III-D to maintain the good performance of the authorized source domain and reduce the generalization of the unauthorized target domains. To fulfill Definition 1, the training objective of the protected model $f_{\theta_{j}}(\cdot)$ can be formulated as a bi-level optimization problem as follows:
$\mathop{\arg\min}\limits_{\theta_{j}}$ $\mathbb{E}_{(x^{+}_{j},y)\sim\mathcal{A}_{j}}\left[\mathcal{L}_{\text{CE}}\left(f_{\theta_{j}}(x^{+}_{j}),y\right)\right]$ (16)
$+\lambda_{3}\sum_{l\in T_{\text{sel}}}\mathbb{E}_{p(x^{+}_{j},x)}\left[\hat{I}\left(f_{\theta_{j}}^{l}(x^{+}_{j});f_{\theta_{j}}^{l}(x)\right)\right]$
$+\lambda_{3}\sum_{l\in T_{\text{sel}}}\mathbb{E}_{p(x^{+}_{j},x^{-}_{j})}\left[\hat{I}\left(f_{\theta_{j}}^{l}(x^{+}_{j});f_{\theta_{j}}^{l}(x^{-}_{j})\right)\right],$
s.t. $\phi^{l}_{1}=\mathop{\arg\min}\limits_{\phi_{1}}{\mathbb{E}_{p(x^{+}_{j},x)}\left[\hat{I}\left(f_{\theta_{j}}^{l}(x^{+}_{j});f_{\theta_{j}}^{l}(x)\right)\right]},$
$\phi^{l}_{2}=\mathop{\arg\min}\limits_{\phi_{2}}{\mathbb{E}_{p(x^{+}_{j},x^{-}_{j})}\left[\hat{I}\left(f_{\theta_{j}}^{l}(x^{+}_{j});f_{\theta_{j}}^{l}(x^{-}_{j})\right)\right]},$
@@
where $f_{\theta_{j}}^{l}(x)$, $f_{\theta_{j}}^{l}(x^{+}_{j})$ and $f_{\theta_{j}}^{l}(x^{-}_{j})$ denote the $l$-th layer latent features of $\mathcal{B}$, $\mathcal{A}_{j}$ and $\mathcal{N}_{j}$ domains, respectively. $\phi^{l}_{1}$ and $\phi^{l}_{2}$ are parameters of the two auxiliary networks used to estimate the $l$-th layer MI between the source domain and two target domains, i.e., $\hat{I}(f_{\theta_{j}}^{l}(x^{+}_{j});f_{\theta_{j}}^{l}(x))$ and $\hat{I}(f_{\theta_{j}}^{l}(x^{+}_{j});f_{\theta_{j}}^{l}(x^{-}_{j}))$. $\lambda_{3}$ is a factor to balance the contributions of different terms.
@@
Unfortunately, it is computationally intensive to iteratively update the two auxiliary networks in tandem with the protected model. Moreover, the optimization of Eq. (16) is difficult to converge. The trained model often exhibits polarized behavior, consistently classifying test images either correctly (satisfying the first term) or incorrectly (satisfying the last two terms), regardless of whether the images are authorized. To address these issues, we design an MoE [37] model $\mathcal{M}_{j}(\cdot)$ with three components: one real expert model $f_{\theta^{R}_{j}}(\cdot)$ that performs well on $\mathcal{A}_{j}$, and two fake expert models $f_{\theta^{F_{1}}_{j}}(\cdot)$ and $f_{\theta^{F_{2}}_{j}}(\cdot)$ to compromise the performance on two unauthorized domains $\mathcal{B}$ and $\mathcal{N}_{j}$, respectively. The real expert model can be fine-tuned from the unprotected benign model $f_{\theta}(\cdot)$ for a few epochs to ensure it achieves the optimal performance on $\mathcal{A}_{j}$ by optimizing
$\dot{\theta}^{R}_{j}=\mathop{\arg\min}{\mathbb{E}_{(x^{+}_{j},y)\sim\mathcal{A}_{j}}\left[\mathcal{L}_{\text{CE}}\left(f_{\theta^{R}_{j}}(x^{+}_{j}),y\right)\right]}.$ (17)
@@
Conversely, the two fake experts are separately optimized by minimizing their MI with the real expert $f_{\theta^{R}_{j}}(\cdot)$ in selected layers as follows:
$\dot{\theta}^{F_{1}}_{j}=\mathop{\arg\min}\sum_{l\in T_{\text{sel}}}\mathbb{E}_{p(x^{+}_{j},x)}\left[\hat{I}\left(f^{l}_{\theta^{R}_{j}}(x^{+}_{j}),f^{l}_{\theta^{F_{1}}_{j}}(x)\right)\right],$ (18)
$\dot{\theta}^{F_{2}}_{j}=\mathop{\arg\min}\sum_{l\in T_{\text{sel}}}\mathbb{E}_{p(x^{+}_{j},x^{-}_{j})}\left[\hat{I}\left(f^{l}_{\theta^{R}_{j}}(x^{+}_{j}),f^{l}_{\theta^{F_{2}}_{j}}(x^{-}_{j})\right)\right].$ (19)
@@
The training objective of Eq. (16) is decomposed into three terms independently optimized by Eqs. (17) (18) and (19) in three separate models. Since the three optimization terms are now disentangled from each other, the training process of the three experts can converge more quickly, thereby reducing the training cost. The trained MoE model $\mathcal{M}_{j}(\cdot)$ can be expressed as follows:
$\mathcal{M}_{j}(x)=\begin{cases}f_{\theta^{R}_{j}}(x),\ \ &\text{if}\ \ x\sim\mathcal{A}_{j},\\ f_{\theta^{F_{1}}_{j}}(x),\ \ &\text{if}\ \ x\sim\mathcal{B},\\ f_{\theta^{F_{2}}_{j}}(x),\ \ &\text{if}\ \ x\sim\mathcal{N}_{j}.\\ \end{cases}$ (20)
@@
We separately feed 1500 paired test images sampled from $\mathcal{W}=\mathcal{A}_{j}\cup\mathcal{B}\cup\mathcal{N}_{j}$ (i.e., 500 samples for each subdomain) and visualize their latent features with t-distributed Stochastic Neighbor Embedding (t-SNE) technique [38]. As shown in Fig. 3, the latent features form three distinct clusters corresponding to their respective subdomains. As expected, the authorized domain shares minimal MI with the two unauthorized domains in the latent space. Consequently, the latent features from different domains exhibit increased separation.
@@
IV-C Model Obfuscation
It is worth noting that the MoE model $\mathcal{M}_{j}(\cdot)$ lacks a gating network to decide which expert should be activated for prediction. Consequently, $\mathcal{M}_{j}(\cdot)$ requires prior knowledge about the domain from which the input image is sampled to activate the correct domain expert, which is impractical in real-world scenarios. Additionally, distributing $\mathcal{M}_{j}(\cdot)$ directly to a user poses a security risk, as adversaries can easily identify the real expert due to its superior classification performance on benign test samples compared to the two fake experts. To address these issues, we obfuscate the three experts by unifying them into a single coherent model using knowledge distillation. We use $\mathcal{M}_{j}(\cdot)$ to train a student model $f_{\theta^{p}_{j}}(\cdot)$ with the secret key $k_{j}$ for distributing to an authorized user. We want $f_{\theta^{p}_{j}}(\cdot)$ to imitate the teacher’s behaviors over $\mathcal{W}$ on not only the fully-connected layer but all selected hidden layers from shallow to deep. To achieve this, we jointly optimize $f_{\theta^{p}_{j}}(\cdot)$ with three knowledge distillation losses $\mathcal{L}_{\text{kl}}$, $\mathcal{L}_{\text{at}}$ and $\mathcal{L}_{\text{crd}}$ such that
$\dot{\theta}^{p}_{j}=\mathop{\arg\min}\left[\mathcal{L}_{\text{kl}}+\lambda_{1}\mathcal{L}_{\text{at}}+\lambda_{2}\mathcal{L}_{\text{crd}}\right],$ (21)
where $\lambda_{1}$ and $\lambda_{2}$ are two positive factors to adjust the scale of $\mathcal{L}_{\text{at}}$ and $\mathcal{L}_{\text{crd}}$, respectively.
$\mathcal{L}_{\text{kl}}=\mathbb{E}_{x\sim\mathcal{W}}\left[\mathcal{L}_{\text{KL}}\left(f_{\theta^{p}_{j}}(x),\mathcal{M}_{j}(x)\right)\right]$ (22)
(2) A 3D feature map $f^{l}\in\mathbb{R}^{{c_{l}}\times{h_{l}}\times{w_{l}}}$ can be flatten along the channel dimension by the attention operator $\mathcal{A}^{\alpha}(\cdot)$:
$\mathcal{A}^{\alpha}(f^{l})=\sum^{c_{l}}_{i=1}\left|f^{l}_{i}\right|^{\alpha},$ (23)
where $\alpha>1$ is the factor that amplifies large neuron activations. Enlarging $\alpha$ places greater emphasis on regions with high output values. $\alpha$ is fixed to 2 according to [39].
The attention transfer loss [40] is computed as:
$\mathcal{L}_{\text{AT}}(f^{l}_{\text{T}},f^{l}_{\text{S}})=\left\|\frac{\mathcal{A}^{\alpha}\left(f^{l}_{\text{T}}\right)}{\left\|\mathcal{A}^{\alpha}\left(f^{l}_{\text{T}}\right)\right\|_{2}}-\frac{\mathcal{A}^{\alpha}\left(f^{l}_{\text{S}}\right)}{\left\|\mathcal{A}^{\alpha}\left(f^{l}_{\text{S}}\right)\right\|_{2}}\right\|_{2},$ (24)
where $f^{l}_{\text{T}}$ and $f^{l}_{\text{S}}$ represent the features extracted from the teacher and student models at the $l$-th layer, respectively.
We also align the neuron attention of $f_{\theta^{p}_{j}}(\cdot)$ and $\mathcal{M}_{j}(\cdot)$ on the selected layers by minimizing
$\mathcal{L}_{\text{at}}=\mathbb{E}_{x\sim\mathcal{W}}\left[\sum_{l\in T_{\text{sel}}}\mathcal{L}_{\text{AT}}\left(f^{l}_{\theta^{p}_{j}}(x),\mathcal{M}^{l}_{j}(x)\right)\right].$ (25)
(3) To ensure that input samples from different domains also share low MI in the latent space in the student model, contrastive representation distillation (CRD) [41] loss is incorporated. Given a tuple of teacher and student representations $(f^{l}_{T}(x_{a}),f^{l}_{S}(x_{b}))$, where $x_{a}$ and $x_{b}$ are two random images sampled from $\mathcal{W}$. It is a positive pair if the two inputs are from the same domain (e.g., $x_{a}\sim\mathcal{A}_{j}$ and $x_{b}\sim\mathcal{A}_{j}$). Conversely, it is a negative pair if the two inputs are sourced from different domains (e.g., $x_{a}\sim\mathcal{A}_{j}$ and $x_{b}\sim\mathcal{B}$). Let $q(f^{l}_{T},f^{l}_{S}|d=1)$ and $q(f^{l}_{T},f^{l}_{S}|d=0)$ denote the joint distributions for positive and negative pairs, respectively. The CRD loss is given by:
$\mathcal{L}_{\text{crd}}=$ $-\sum_{l\in T_{\text{sel}}}\mathbb{E}_{q(f^{l}_{T},f^{l}_{S}|d=1)}\left[\text{log}\left(h^{l}(f^{l}_{T},f^{l}_{S})\right)\right]$ (26)
$-N_{\text{neg}}\sum_{l\in T_{\text{sel}}}\mathbb{E}_{q(f^{l}_{T},f^{l}_{S}|d=0)}\left[\text{log}\left(1-h^{l}(f^{l}_{T},f^{l}_{S})\right)\right],$
where $h^{l}(\cdot,\cdot)$ denotes a critic to determine if the two representations are positively paired, and whose parameters can be optimized in parallel with $\mathcal{L}_{\text{crd}}$. We provide one positive pair for every $N_{\text{neg}}$ negative pairs during training, with the aim to pull the positive pairs closer and push the negative pairs apart.
The protected model $f_{\theta^{p}_{j}}(\cdot)$ has similar performance as the MoE model $\mathcal{M}_{j}(\cdot)$, but the functionalities and hidden-layer behaviors of the three experts have been indistinguishably blent into a coherent network, making it infeasible for the adversary to extract the real expert model. The algorithm for generating the $M$ encoded models $\{f_{\theta^{p}_{j}}(\cdot),k_{j}\}_{j=1}^{M}$ is depicted in Algorithm 1.
Input: Benign training dataset $D_{\text{tr}}$; unprotected model $f_{\theta}(\cdot)$; SteganoGAN encoder $\mathcal{E}(\cdot,\cdot)$; $M$ unique user keys $\mathcal{K}=\{k_{1},k_{2},\cdots,k_{M}\}$; training epochs $\{E_{1},E_{2},E_{3}\}$; learning rates $\{\eta_{1},\eta_{2},\eta_{3}\}$; set of selected hidden layers $T_{\text{sel}}$.
@@ IV-C Model Obfuscation
Output: $M$ specifically-encoded models $\{f_{\theta^{p}_{1}}(\cdot),f_{\theta^{p}_{2}}(\cdot),\cdots,f_{\theta^{p}_{M}}(\cdot)\}$.
@@
6       $D^{+}_{\text{tr},j}\leftarrow\{(\mathcal{E}(x_{i},k_{j}),y_{i})\}^{N_{\text{tr}}}_{i=1}$

### v2

@@ III-C SteganoGAN
SteganoGAN [27] is a lightweight steganographic network for key encoding and decoding. It consists of three sub-networks: an encoder $\mathcal{E}$ which imperceptibly embeds binary messages into input images; a decoder $\mathcal{D}$ which extracts the embedded messages from encoded images; and a critic $\mathcal{C}$ reduces distortion in the encoded images. SteganoGAN is employed in IDEA to hide and retrieve secret message imperceptibly into the input image due to its simple architecture, which reduces the computational overhead for key embedding and extraction.
@@ III-D Domain Expected Risk Bound
$\mathcal{R}_{\mathcal{T}}(f)\leq\mathcal{R}_{\mathcal{S}}(f)-4I(z;\hat{z})+4H(Y)+\frac{1}{2}d_{\mathcal{H}\Delta\mathcal{H}}\left(p(z),p(\hat{z})\right),$ (7)
@@
The unknown conditional distribution $p(\hat{z}|z)$ can be approximated by the variational distribution $q_{\phi}(\hat{z}|z)$ with an auxiliary neural network parameterized with $\phi$ [44]. By feeding samples $\{(z_{i},\hat{z}_{i})\}_{i=1}^{N_{\text{est}}}$ into the auxiliary network, the MI between two feature distributions can be estimated as:
@@
Previous works [43, 35] focused on minimizing the MI between the latent features output by the penultimate layer to reduce the generalization of $\mathcal{T}$. We observe that the source and target domain samples share minimal information at the penultimate layer. Their latent features in the earlier layers are fairly similar, which can leak vital information about the protected model’s functionality. To avoid this vulnerability, we jointly minimize the MI between $\mathcal{S}$ and $\mathcal{T}$ on 4 selected hidden layers (i.e., $T_{\text{sel}}=\{l_{1},l_{2},l_{3},l_{4}\}$) that are evenly distributed from shallow to deep within the model.
@@ IV-A Key Embedding
Fig. 2 illustrates the process of key encoding and decoding. Each user is assigned a unique key ${k}\in{\{0,1\}}^{{r}\times{r}}$, a binary square matrix where the first half encodes the owner’s identity and the second half encodes the user’s identity.
SteganoGAN requires the input image and the key to have matching dimensions in height and width. However, empirical observations indicate that directly matching the key size to the image size is inefficient. For high-resolution datasets like ImageNet, this approach results in an enormous key space of $2^{224\times 224}$, which is unnecessary given the limited number of authorized users. Additionally, such a large key space hinders the protected model’s ability to distinguish the correct key from incorrect ones, reducing the uniqueness of active authorization. To address this, the key size $r$ is fixed to values like 16 or 32, and multiple non-overlapping blocks of $k$ are concatenated to create an expanded key to match the image size. The encoded image is then generated by feeding the input image and the expanded key into the encoder simultaneously.
During the verification stage, the encoded image is passed through the decoder and split into multiple non-overlapping blocks of size ${r}\times{r}$. A majority voting process is then applied element-wise across the blocks to reconstruct the decoded key $\hat{k}$, which improves the decoding accuracy.
The encoder is provided to authorized users for embedding their unique keys into test images, while the decoder is kept private by the model owner to extract the embedded keys from encoded images. The perturbations introduced by the encoder are visually imperceptible and sample-specific, which implies that the perturbations vary across images even for the same embedded key.
Figure 1: Overall pipeline of IDEA, a solution to the inverse problem of domain adaptation for key-based active authorization. The owner trains three experts: one real expert, which performs well on the authorized domain $\mathcal{A}$, and two fake experts, which behave like untrained models on the benign domain $\mathcal{B}$ and the noise domain $\mathcal{N}$, respectively. These three experts are seamlessly integrated into a unified model to enable active control while preventing the real expert’s functionality from being leaked. To verify a licensee, the owner can extract a specific user key from encoded images provided by the user. Furthermore, the owner can conduct black-box queries on a suspected model to confirm the ownership and trace the dishonest user responsible for redeployment or infringement.
Figure 2: Key embedding into a cover image by the pretrained SteganoGAN encoder and key extraction from the stego image by the model owner with the corresponding pretrained SteganoGAN decoder (private). The SteganoGAN encoder and decoder are pretrained on the original training dataset.
@@ IV-B Experts Training
Let $\mathcal{A}$ be the source domain, and $\mathcal{B}$ and $\mathcal{N}$ be two different target domains. An inverse problem of domain adaption can be solved by optimizing the domain expected risk bound presented in Sec. III-D to maintain the good performance of the authorized source domain and reduce the generalization of the unauthorized target domains. To fulfill Definition 1, the training objective of the protected model $f^{\prime}$ can be formulated as a bi-level optimization problem as follows:
$\mathop{\min}$ $\mathbb{E}_{(x^{+},y)\sim\mathcal{A}}\left[\mathcal{L}_{\text{CE}}\left(f^{\prime}(x^{+}),y\right)\right]$ (10)
$+\lambda_{1}\sum_{l\in T_{\text{sel}}}\left[\hat{I}\left(f^{\prime}_{l}(x^{+});f^{\prime}_{l}(x)\right)\right]$
$+\lambda_{1}\sum_{l\in T_{\text{sel}}}\left[\hat{I}\left(f^{\prime}_{l}(x^{+});f^{\prime}_{l}(x^{-})\right)\right],$
s.t. $\phi_{l,1}=\mathop{\arg\min}\limits_{\phi_{l,1}}\ {\hat{I}\left(f^{\prime}_{l}(x^{+});f^{\prime}_{l}(x)\right)},$
$\phi_{l,2}=\mathop{\arg\min}\limits_{\phi_{l,2}}\ {\hat{I}\left(f^{\prime}_{l}(x^{+});f^{\prime}_{l}(x^{-})\right)},$
@@
where $f^{\prime}_{l}(x)$, $f^{\prime}_{l}(x^{+})$ and $f^{\prime}_{l}(x^{-})$ denote the $l$-th layer latent features of $\mathcal{B}$, $\mathcal{A}$ and $\mathcal{N}$ domains, respectively. $\hat{I}(\cdot;\cdot)$ denotes the estimated MI between two domains. $\phi_{l,1}$ and $\phi_{l,2}$ are parameters of the two auxiliary networks used to estimate the $l$-th layer MI between the source domain and two target domains. $\lambda_{1}$ is a factor to balance the contributions of different terms.
@@
Unfortunately, it is computationally intensive to iteratively update the two auxiliary networks in tandem with the protected model. Moreover, the optimization of Eq. (10) is difficult to converge. The trained model often exhibits polarized behavior, consistently classifying test images either correctly (satisfying the first term) or incorrectly (satisfying the last two terms), regardless of whether the images are authorized.
To address these issues, we design an MoE [45] model $\mathcal{M}$ with three expert models: one real expert that performs well on $\mathcal{A}$, and two fake expert models that deliver random guessing performance on two unauthorized domains $\mathcal{B}$ and $\mathcal{N}$, respectively. The three experts have the same architecture as the unprotected model $f$.
The real expert $f_{\text{real}}$ is fine-tuned from $f$ for a few epochs to ensure that it achieves the optimal performance on $\mathcal{A}$ by optimizing
$\mathop{\min}\ {\mathbb{E}_{(x^{+},y)\sim\mathcal{A}}\left[\mathcal{L}_{\text{CE}}\left(f_{\text{real}}(x^{+}),y\right)\right]}.$ (11)
@@
Conversely, the two fake experts, $f_{\text{fake1}}$ and $f_{\text{fake2}}$, are separately optimized by minimizing the MI of their latent representations with those of the real expert in selected layers as:
$\mathop{\min}\sum_{l\in T_{\text{sel}}}\hat{I}\left(f_{\text{real},l}(x^{+});f_{\text{fake1},l}(x)\right),$ (12)
$\mathop{\min}\sum_{l\in T_{\text{sel}}}\hat{I}\left(f_{\text{real},l}(x^{+});f_{\text{fake2},l}(x^{-})\right),$ (13)
where $f_{\text{real},l}(x^{+})$, $f_{\text{fake1},l}(x)$, and $f_{\text{fake2},l}(x^{-})$ represent the $l$-th layer latent representations of $f_{\text{real}}$ on an authorized sample, $f_{\text{fake1}}$ on a benign sample, and $f_{\text{fake2}}$ on a noise sample, respectively.
@@
The training objective of Eq. (10) is decomposed into three terms independently optimized by Eqs. (11), (12) and (13) in three separate models. Since the three optimization terms are now disentangled from each other, the training process of the three experts can converge more quickly, thereby reducing the training cost. The MoE model behaves as the real expert when fed with authorized images, but produces incorrect predictions, acting as one of the two fake experts when encountering benign or noise images.
@@
We separately feed 1500 paired test images sampled from $\mathcal{W}=\mathcal{A}\cup\mathcal{B}\cup\mathcal{N}$ (i.e., 500 samples for each subdomain) and visualize their latent features with t-distributed Stochastic Neighbor Embedding (t-SNE) technique [46]. As shown in Fig. 3, the latent features form three distinct clusters corresponding to their respective subdomains. As expected, the authorized domain shares minimal MI with the two unauthorized domains in the latent space. Consequently, the latent features from different domains exhibit increased separation.
@@
Input: Benign training dataset $D_{\text{tr}}$; unprotected model $f$; SteganoGAN encoder $\mathcal{E}$; $M$ unique user keys $\mathcal{K}=\{k_{1},k_{2},\cdots,k_{M}\}$; training epochs $\{E_{1},E_{2},E_{3}\}$; learning rates $\{\eta_{1},\eta_{2},\eta_{3}\}$; set of selected hidden layers $T_{\text{sel}}$.
@@
Output: $M$ model-key pairs.
@@
6       $D^{+}_{\text{tr},j}\leftarrow\{(\mathcal{E}(x_{i},k),y_{i})\}^{N_{\text{tr}}}_{i=1}$